please help me solve
find two consecutive whole numbers such that 5 times the smaller number added to 3 times the greater number makes 59

Answer by ankor@dixie-net.com(22740)

find two consecutive whole numbers
x, x+1
such that 5 times the smaller number added to 3 times the greater number makes 59
5x + 3(x+1) = 59
5x + 3x + 3 = 59
8x = 59 - 3
8x = 56
x = 56/8
x = 7
:
numbers are 7 and 8, you can confirm this in the original equation
5(7) + 3(8)